# A monolithic first–order BSSNOK formulation of the Einstein–Euler equations and its solution with path-conservative finite difference CWENO schemes

### v1

@@ I Introduction
The two pillars of any successful numerical simulation are represented by an appropriate formulation of the equations that needs to be solved and by the availability of an accurate and robust numerical scheme. Concerning the formulation of the Einstein equations, the vast majority of numerical relativity codes that are currently adopted to compute gravitational waveforms is based on the celebrated (second–order in space) BSSNOK formulation [13]. In spite of a few deprecated defects, the most annoying of which is perhaps the failure in keeping a simple Kerr black hole stationary over a long timescale, this formulation has shown a spectacular success in reproducing the merger of binary systems, which was the real busillis of numerical relativity (NR). Precisely because the second–order implementation of the BSSNOK formulation has been so successful in modeling the gravitational wave signals from binary systems, the scientific research on first–order formulations has been somewhat left aside during the last decade or more. As a matter of fact, most of the analysis about first–order formulations of the Einstein equations belong to the beginning of the NR era [14, 15, 16, 17, 18, 19, 20, 21, 22], with only few exceptions in relatively recent years [23, 24, 25, 26, 27].
@@
Concerning the numerical scheme, on the other hand, it is fair to say that, in spite of considerable efforts carried out in the last decade for adopting innovative numerical schemes in the solution of the Einstein equations, most of the existing numerical relativity codes still rely on traditional central finite differencing with Kreiss-Oliger dissipation, in combination with Runge–Kutta time integration. Of course there are exceptions to this trend. For example, there is a active field of research to extend discontinuous Galerkin (DG) schemes to full general relativity [28, 29, 30, 24, 31, 32, 33, 27]; or to develop spectral methods both for the initial value problem and for the time evolution of the Einstein equations [34, 35, 36, 37]; or to go beyond Runge–Kutta schemes for the time integration by migrating to so–called ADER schemes [38, 39, 40, 41, 42, 43]. In spite of these progresses, if we list the most popular numerical relativity codes that perform actual calculations of waveforms from binary mergers (Einstein–Toolkit [^2] http://einsteintoolkit.org/ [44], LazEv [^3] https://ccrg.rit.edu/content/software/lazev [45, 46], BAM [^4] http://data.cardiffgravity.org/bam-catalogue/ [47, 48, 49], GRChombo  [^5] https://www.grchombo.org/ [50, 51], AMReX [^6] https://amrex-codes.github.io/amrex/ [52], Nmesh [33], SACRA  [^7] https://www.aei.mpg.de/1095648/SACRA-description[53, 54], SpEC [^8] https://www.black-holes.org/code/SpEC.html [55, 35, 56], SPHINCS_BSSN [57], SENR/NRPy [58]) we observe that, apart from Nmesh and SpEC, all of them still uses finite difference schemes for the Einstein sector of the spacetime evolution.
@@
Now, if we combine our considerations about the formulation of the Einstein equations, which rests on the second–order BSSNOK system, and the choice of a suitable numerical scheme, which rests mainly on traditional finite differencing, we see that the lack of a reliable first–order formulation represents a serious obstacle for the application of the whole class of advanced numerical schemes of Godunov type proposed in the last three decades (or more) for the solution of first–order hyperbolic PDE systems. As an additional complication, when the Einstein equations are cast in first–order form, they naturally come up with non–conservative terms and curl-type involutions, which may give the (erroneous) impression of being untractable by Godunov–type methods. On the contrary, since all the characteristic fields arising from the Einstein equations are linearly degenerate, discontinuities cannot develop, and therefore non–conservative terms in the PDE system can be taken into account in a relatively easy way. With the benefit of hindsight, we could say that insisting in finding first–order and conservative formulations of the Einstein equations has been counterproductive.
@@
In light of the above comments, assuming that finite difference numerical schemes are likely to remain the preferred choice by the NR community for a certain amount of time onward, partly due to their simplicity and certainly for their efficiency in terms of CPU time performances, in this paper we resume a central WENO (CWENO) finite difference numerical scheme, proposed in a different context originally by [59, 60, 61], that may contribute to a substantial improvement with respect to the present standard in a few specific aspects:
@@
At this point we would like to emphasize that, to the best knowledge of the authors, this is the first time that a fully coupled first order hyperbolic BSSNOK formulation of the Einstein-Euler equations is proposed and which is solved in a totally monolithic manner within one and the same computer code using exactly the same numerical scheme for both matter and spacetime.
@@ II The BSSNOK system
II.1 The standard second order version
The vast majority of numerical codes solving the Einstein–Euler equations within the $3+1$ formalism adopt the so–called BSSNOK formulation, which, due to its popularity, does not require an extended discussion and it can be briefly described as follows (for more details see  [73, 13, 74, 75]). The spatial metric $\gamma_{ij}$ of the spacelike hypersurfaces is rescaled conformally according to
@@ II.1 The standard second order version
$(\partial_{t}-{\cal L}_{\beta})~{}\tilde{\gamma}_{ij}$ $=$ $-{2\,\alpha}\,\tilde{A}_{ij}\,,$ (3)
$(\partial_{t}-{\cal L}_{\beta})~{}\phi$ $=$ $-\frac{1}{6}\alpha\,K\,,$ (4)
$(\partial_{t}-{\cal L}_{\beta})~{}\tilde{A}_{ij}$ $=$ $e^{-4\phi}\bigg{[}-\nabla_{i}\nabla_{j}\alpha+\alpha\,R_{ij}-8\pi\alpha S_{ij}\,\bigg{]}^{TF}+\alpha\left(K\,\tilde{A}_{ij}-2\tilde{A}_{im}\tilde{A}^{m}_{j}\right)\,,$ (5)
$(\partial_{t}-{\cal L}_{\beta})~{}K$ $=$ $-\nabla_{i}\nabla^{i}\alpha+\alpha\left(\tilde{A}_{ij}\tilde{A}^{ij}+\frac{1}{3}K^{2}+4\pi(E+S)\right)\,,$ (6)
$(\partial_{t}-{\cal L}_{\beta})~{}\tilde{\Gamma}^{i}$ $=$ $\tilde{\gamma}^{jk}\partial_{j}\partial_{k}\beta^{i}+\frac{1}{3}\tilde{\gamma}^{ij}\partial_{j}\partial_{k}\beta^{k}-2\tilde{A}^{ij}\partial_{j}\alpha$
@@
and these quantities will be monitored during a numerical simulation. In particular, we will check the $L_{2}$ errors of the above constraints, computed as
$L_{2}=\int_{\Omega}\,\epsilon^{2}\,\sqrt{\gamma}\,d^{3}x\,\hskip 28.45274pt\textrm{with}\hskip 28.45274pt\epsilon=H,M^{i}\,,$ (10)
where no normalization is performed with respect to the volume of integration.
@@ II.2 The new first–order version
A promising first–order formulation of BSSNOK was proposed by [23], though it seems that traditional second order formulations are still the preferred choice in large scale simulations of astrophysical sources. Here, as already done in the first–order CCZ4 formulation of [24], we introduce 30 auxiliary variables containing first derivatives of the metric terms, namely
@@
that can be used in the first order version of BSSNOK. With the aid of (11), the system (3)–(II.1), augmented by the matter part and by the gauge conditions, can be written as a monolithic first–order BSSNOK formulation of the Einstein-Euler system, i.e.,
@@
$\partial_{t}A_{i}-{\beta^{k}\partial_{k}A_{i}}+\alpha g(\alpha)\left(\partial_{i}K-\partial_{i}K_{0}\right)=-\alpha A_{i}\left(K-K_{0}\right)\left(g(\alpha)+\alpha g^{\prime}(\alpha)\right)+B_{i}^{\,\,k}~{}A_{k}$ (25)
@@
  • The coefficient $\mu$ in Eq. (26) can be used to modulate the insertion of the ordering constraints into the system, specifically the red terms $\gamma^{ij}\gamma^{nl}\left(\partial_{k}D_{ljn}-\partial_{l}D_{kjn}\right)$, which are zero at the continuous level due to the symmetry of second derivatives, see Eq. (12).
  • The extra red term proportional to $\left[\tilde{\gamma}^{jk}\partial_{i}\tilde{A}_{jk}+\tilde{A}_{jk}\partial_{i}\tilde{\gamma}^{jk}\right]$ on the right hand side of Eq. (II.2), is added to take into account that $\tilde{A}_{jk}$ is trace free, see also [24], and are crucial to obtain a strongly hyperbolic system.
@@
The definition of the matter quantities $D$, $E$, $S_{i}$, $S_{ij}$ for a non–dissipative fluid in Eq. (14)–(16) is pretty standard and it can be found, for instance, in [75]. Throughout this paper we are assuming a simple ideal gas equation of state $p=\rho\epsilon(\gamma-1)$, where $\epsilon$ is the specific internal energy, while $\gamma$ is the adiabatic index. Of course, present day NR simulations adopt much more realistic equations of state, especially for neutron star binaries (see, among the others, [76, 77, 78]), but this is not the focus of this work, and we have therefore set up the simplest physical conditions. For the sake of completeness we recall the definitions of the conservative variables and of the stress tensor in terms of the primitive (physical) variables:
@@ III.1 Central WENO reconstruction
Central WENO (CWENO) schemes were proposed more than 20 years ago in a series of papers by [59, 60, 61]. Here we provide a step-by-step presentation of this approach, showing how it can be practically implemented within the framework of conservative WENO finite difference methods [65, 66, 67]. We consider for simplicity the $x$-direction, while the extension to a 3D implementation can be easily obtained after repeating the same reconstruction also in the $y$ and $z$ directions. We assume the discrete solution of Eq. (59) below to be given by the point values $\mathbf{u}_{i,j,k}=\mathbf{u}(x_{i},y_{j},z_{k})$, with the points located in the centers of the logically equidistant control volumes, i.e. $x_{i}=\frac{1}{2}(x_{i-\frac{1}{2}}+x_{i+\frac{1}{2}})$ and $\Delta x=x_{i+\frac{1}{2}}-x_{i-\frac{1}{2}}$. More precisely, we have
@@
    In each direction, there is a total stencil $\mathcal{S}_{opt}$ given by $\mathcal{S}_{opt}=\left\{i-r,\cdots,i-1,i,i+1,\cdots,i+r\right\}$ with $r=N/2$ the half stencil size and $N$ the degree of the polynomial to be reconstructed on such a stencil. In addition, there are a number of sub-stencils, which we list below according to the value of $N$:
@@
    A generic polynomial of degree $M$ on any stencil $k$ is expressed in terms of a set of polynomial basis functions $\psi_{m}(x)$ and associated degrees of freedom $\hat{u}_{k}^{m}$ as follows:
@@
    Throughout this paper we will use the simple Taylor monomials $\psi_{m}(x)=x^{m}$ as basis functions. The polynomials $P^{M}(x)$ are obtained by reconstruction, i.e. by interpreting the point values $\mathbf{u}_{i,j,k}$ of the finite difference scheme as cell averages, see [66]. The associated reconstruction equations (which are not interpolation equations, see [66, 67]) for a generic polynomial of degree $M$ computed on an associated stencil $\mathcal{S}_{k}$ read
@@
    Consequently, from (50) and (51) we now compute the auxiliary polynomial as
@@
    Finally, a non-linear WENO reconstruction is performed in terms of the polynomial $P^{N}_{0}(x)$ and of the lower degree polynomials built over the existing sub-stencils. In particular, we compute for all polynomials of degree $M(k)$ the nonlinear WENO oscillation indicators as follows
@@
so that the corresponding spatial derivative can be obtained as
@@ III.2 CWENO finite difference discretization of the Einstein-Euler system
The full matter + Einstein equations (14)–(II.2) form a first–order PDE system in which the matter sector enters as a purely conservative sub–system, while the Einstein sector is purely non–conservative. On the overall it can be written as
@@
where $\boldsymbol{u}$ is the state vector, composed of 63 dynamical variables [^9] More specifically, 5 for the matter part, 11 for the lapse, the shift vector, the metric components and the scalar $\phi$, 7 for $\tilde{A}_{ij}$ and the scalar $K$, 3 for $A_{i}$, 9 for $B_{i}^{\,\,j}$, 18 for $D_{ijk}$, 3 for $P_{i}$, 1 for $K_{0}$, 3 for $b^{i}$ and 3 for $\tilde{\Gamma}^{i}$., while $\boldsymbol{F}=\left(\boldsymbol{f}^{x},\boldsymbol{f}^{y},\boldsymbol{f}^{z}\right)^{T}$ is the flux vector. A finite difference WENO scheme for the vacuum Einstein equations, namely with $T^{\mu\nu}=0$ and which can successfully cope with the presence of non–conservative terms, has been recently proposed by [64]. Here we follow a similar approach, with a few modifications. Still in semi-discrete form, a path-conservative finite difference scheme for the discretization of (59), reads
@@
Note that one only needs the eigenvalues, not the matrix itself. Eq. (61) and (62) should be regarded as four–state flux formulas, where each of the four entries $\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}$, $\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k}$, $\boldsymbol{f}^{x,-}_{i+\frac{1}{2},j,k}$ and $\boldsymbol{f}^{x,+}_{i+\frac{1}{2},j,k}$ are reconstructed via the CWENO strategy, as described in Sect. III.1. The whole procedure must of course be repeated for the calculation of $\boldsymbol{f}^{y}_{i,j\pm\frac{1}{2},k}$ and $\boldsymbol{f}^{z}_{i,j,k\pm\frac{1}{2}}$. According to [72, 64] the three terms highlighted in red in Eq. (60) are equivalent to the jump terms in path-conservative finite volume schemes [80, 81, 82, 83] and they are given as follows:
  • Rusanov–type dissipation:
@@
The computation of the terms $\tilde{\boldsymbol{B}}^{x}_{i+\frac{1}{2},j,k}$ in (67) and (III.2) is performed through a simple midpoint rule as
@@
where $\xi_{i}^{c}$ sets the border of the inner uniform grid, while the coefficients $a$, $b$, $c$ and $d$ are selected in such a way to obtain a smooth matching of the cubic with the inner linear branch. The same mapping is applied along all coordinate directions.
@@
For stationary (or quasi stationary) problems, our numerical scheme can easily be made well–balanced, by simply subtracting a discrete version of the equilibrium solution from the discretized time-dependent PDE system. In this way, an initial equilibrium can be maintained stationary up to machine precision virtually for ever. The interested reader is referred to [85, 86, 27], where he/she can find all the practical details for the simple and straightforward implementation of a well–balanced scheme in a PDE system of mixed conservative///non–conservative form. For more theoretical background on well-balanced numerical methods for hyperbolic PDE, see [87, 88, 89, 90, 91, 80, 81, 82] and references therein.

### v2

@@ I Introduction
The two main pillars of any successful numerical simulation are represented by an appropriate mathematical formulation of the equations that need to be solved and by the availability of an accurate and robust numerical scheme.
Concerning the mathematical formulation of the Einstein field equations, although many of them have been proposed over the years, both as first–order or second–order PDEs systems, two of them are the most adopted in numerical relativity codes today: $i)$ the generalized harmonic formulation (GH), which is usually written as a first order (in space derivatives) PDE system [13], with very important scientific results, including the first stable numerical simulation of a binary black hole merger presented in [14], and $ii)$ the second–order (in space derivatives) BSSNOK formulation (with its variants represented by CCZ4/Z4c) within the 3+1 splitting of spacetime [15, 16, 17]. In this article we are concerned only with the latter approach. The scientific research on first–order formulations of the BSSNOK system has been somewhat left aside during the last decade or more, with only a few exceptions in relatively recent years [18, 19, 20, 21].
@@
Concerning the numerical scheme, on the other hand, it is fair to say that, in spite of considerable efforts carried out in the last decade for adopting innovative numerical schemes in the solution of the Einstein equations, most of the existing numerical relativity codes still rely on rather traditional central finite differencing with artificial dissipation of the Kreiss-Oliger type, in combination with Runge–Kutta time integration. Of course there are exceptions to this trend. For example, there is a active field of research to extend discontinuous Galerkin (DG) schemes to full general relativity [22, 23, 24, 20, 25, 26, 27, 28, 29, 30]; or to develop spectral methods both for the initial value problem and for the time evolution of the Einstein equations [31, 32, 33, 34, 35, 36, 37]; or to go beyond Runge–Kutta schemes for the time integration by migrating to so–called ADER schemes [38, 39, 40, 41, 42, 43]. In spite of these progresses, if we list the most popular numerical relativity codes that perform actual calculations of waveforms from binary mergers (Einstein–Toolkit [^2] http://einsteintoolkit.org/ [44], LazEv [^3] https://ccrg.rit.edu/content/software/lazev [45, 46], BAM [^4] http://data.cardiffgravity.org/bam-catalogue/ [47, 48, 49], GRChombo [^5] https://www.grchombo.org/ [50, 51], AMReX [^6] https://amrex-codes.github.io/amrex/ [52], Nmesh [27], SACRA [^7] https://www.aei.mpg.de/1095648/SACRA-description[53, 54], SpEC [^8] https://www.black-holes.org/code/SpEC.html [55, 33, 56], SPHINCS_BSSN [57], SENR/NRPy [58], MHDueT [^9] http://mhduet.liu.edu/[59, 60]), GR-Dendro [^10] https://paralab.github.io/Dendro-GR/ [61, 62]) we observe that, apart from Nmesh, SpEC and GR-Dendro, all of them still use finite difference schemes for the Einstein sector of the spacetime evolution.
@@
Now, if we combine our considerations about the formulation of the Einstein equations, which rests on the second–order BSSNOK system, and the choice of a suitable numerical scheme, which rests mainly on traditional finite differencing, we see that the lack of a reliable first–order formulation represents a serious obstacle for the application of the whole class of advanced numerical schemes of Godunov-type proposed in the last three decades (or more) for the solution of nonlinear first–order hyperbolic PDE systems. As an additional complication, when the Einstein equations are cast in first–order form, they naturally come up with non–conservative terms and curl-type involutions, which may give the (erroneous) impression of being untractable by Godunov–type methods [^11] For an alternative mixed elliptic-hyperbolic formulation see also [63, 64].. On the contrary, since all the characteristic fields arising from the Einstein equations are linearly degenerate, discontinuities cannot develop, and therefore non–conservative terms in the PDE system can be taken into account in a relatively easy way.
@@
In light of the above comments, assuming that finite difference numerical schemes are likely to remain the preferred choice by the NR community for a certain amount of time onward, due to their simplicity, in this paper we resume a central WENO (CWENO) finite difference numerical scheme, proposed in a different context originally by [65, 66, 67], that may contribute to a substantial improvement with respect to the present standard in a few specific aspects:
@@ II The BSSNOK system
II.1 The standard second order formulation
Many numerical codes solving the Einstein–Euler equations within the $3+1$ formalism adopt the so–called second order BSSNOK formulation, which, due to its popularity, does not require an extended discussion and it can be briefly described as follows (for more details see  [16, 17, 79, 80]). The spatial metric $\gamma_{ij}$ of the spacelike hypersurfaces is rescaled conformally according to
@@ II.1 The standard second order formulation
$(\partial_{t}-{\cal L}_{\beta})\leavevmode\nobreak\ \tilde{\gamma}_{ij}$ $=$ $-{2\,\alpha}\,\tilde{A}_{ij}\,,$ (3)
$(\partial_{t}-{\cal L}_{\beta})\leavevmode\nobreak\ \phi$ $=$ $-\frac{1}{6}\alpha\,K\,,$ (4)
$(\partial_{t}-{\cal L}_{\beta})\leavevmode\nobreak\ \tilde{A}_{ij}$ $=$ $e^{-4\phi}\bigg{[}-\nabla_{i}\nabla_{j}\alpha+\alpha\,R_{ij}-8\pi\alpha S_{ij}\,\bigg{]}^{TF}+\alpha\left(K\,\tilde{A}_{ij}-2\tilde{A}_{im}\tilde{A}^{m}_{j}\right)\,,$ (5)
$(\partial_{t}-{\cal L}_{\beta})\leavevmode\nobreak\ K$ $=$ $-\nabla_{i}\nabla^{i}\alpha+\alpha\left(\tilde{A}_{ij}\tilde{A}^{ij}+\frac{1}{3}K^{2}+4\pi(E+S)\right)\,,$ (6)
$(\partial_{t}-{\cal L}_{\beta})\leavevmode\nobreak\ \tilde{\Gamma}^{i}$ $=$ $\tilde{\gamma}^{jk}\partial_{j}\partial_{k}\beta^{i}+\frac{1}{3}\tilde{\gamma}^{ij}\partial_{j}\partial_{k}\beta^{k}-2\tilde{A}^{ij}\partial_{j}\alpha$
@@
and these quantities will be monitored during a numerical simulation. In particular, we will check the normalized $L^{2}$ errors of the above constraints, computed for a generic quantity $q$ as
$\langle L_{q}^{2}\rangle=\sqrt{\frac{\int_{\Omega}\,q^{2}\,d^{3}x}{\int_{\Omega}\,d^{3}x}}\,\hskip 28.45274pt\textrm{with}\hskip 28.45274ptq\in\left\{H,\,M^{i}\right\}\,.$ (10)
@@ II.2 The new first–order version
A promising first–order formulation of BSSNOK was proposed by [19], though it seems that second–order BSSNOK formulations are still the most popular choice in large scale simulations of astrophysical sources. Here, as already done in the first–order CCZ4 formulation of [20], we introduce 30 auxiliary variables containing first derivatives of the metric terms, namely
@@
that can be used in the first order version of BSSNOK. In this paper we have found it not necessary to apply specific techniques to ensure a curl-free evolution, as done, for instance, in [21] for the first order CCZ4 equations. On the contrary, we have found it crucial, for strong hyperbolicity, to insert the curl-free term proportional to $\mu$ on the left hand side of Eq. (26). This is also discussed in the list of comments below Eq. (43). With the aid of (11), the system (3)–(II.1), augmented by the matter part and by the gauge conditions, can be written as a monolithic first–order BSSNOK formulation of the Einstein-Euler system, i.e.,
@@
$\partial_{t}A_{i}-{\beta^{k}\partial_{k}A_{i}}+\alpha g(\alpha)\left(\partial_{i}K-\partial_{i}K_{0}\right)=-\alpha A_{i}\left(K-K_{0}\right)\left(g(\alpha)+\alpha g^{\prime}(\alpha)\right)+B_{i}^{\,\,k}\leavevmode\nobreak\ A_{k}$ (25)
@@
  • The coefficient $\mu$ in Eq. (26) can be used to modulate the insertion of the ordering constraints into the system, specifically the terms $\gamma^{ij}\gamma^{nl}\left(\partial_{k}D_{ljn}-\partial_{l}D_{kjn}\right)$, which are zero at the continuous level due to the symmetry of second derivatives, see Eq. (12). These terms are crucial to obtain a strongly hyperbolic system.
  • The extra term proportional to $\left[\tilde{\gamma}^{jk}\partial_{i}\tilde{A}_{jk}+\tilde{A}_{jk}\partial_{i}\tilde{\gamma}^{jk}\right]$ on the right hand side of Eq. (II.2), is added to take into account that $\tilde{A}_{jk}$ is trace free, see also [20], and is crucial to obtain a strongly hyperbolic system.
@@
The definition of the matter quantities $D$, $E$, $S_{i}$, $S_{ij}$ for a non–dissipative fluid in Eq. (14)–(16) is standard nowadays and can be found, for instance, in [80]. Throughout this paper we are assuming a simple ideal gas equation of state $p=\rho\epsilon(\gamma-1)$, where $\epsilon$ is the specific internal energy, while $\gamma$ is the adiabatic index. Of course, present day NR simulations adopt much more realistic equations of state, especially for neutron star binaries (see, among the others, [81, 82, 83]), but this is not the focus of this work, and we have therefore set up the simplest physical conditions. For the sake of completeness we recall the definitions of the conservative variables and of the stress tensor in terms of the primitive (physical) variables:
@@ III.1 Central WENO reconstruction
Central WENO (CWENO) schemes were introduced more than 20 years ago in a series of papers by [65, 66, 67]. Here we provide a step-by-step presentation of this approach, showing how it can be practically implemented within the framework of conservative WENO finite difference methods [71, 72, 73, 70]. We consider for simplicity the $x$-direction, while the extension to a 3D implementation can be easily obtained after repeating the same reconstruction also in the $y$ and $z$ directions. We assume the discrete solution of Eq. (59) below to be given by the point values $\mathbf{u}_{i,j,k}=\mathbf{u}(x_{i},y_{j},z_{k})$, with the points located in the centers of the logically equidistant control volumes, i.e. $x_{i}=\frac{1}{2}(x_{i-\frac{1}{2}}+x_{i+\frac{1}{2}})$ and $\Delta x=x_{i+\frac{1}{2}}-x_{i-\frac{1}{2}}$. More precisely, we have
@@
    In each direction, there is a large stencil $\mathcal{S}_{opt}$ composed of $N+1$ points and given by $\mathcal{S}_{opt}=\left\{i-r,\cdots,i-1,i,i+1,\cdots,i+r\right\}$ with $r=N/2$ the half stencil size and $N$ the degree of the polynomial to be reconstructed on such a stencil. In addition, there are a number of sub-stencils for the construction of lower degree polynomials, which we list below according to the value of $N$:
@@
    A generic polynomial of degree $M=M(k)$ on any stencil $\mathcal{S}_{k}$ is expressed in terms of a set of polynomial basis functions $\psi_{m}(x)$ and associated degrees of freedom $\hat{u}_{k}^{m}$ as follows:
@@
    Throughout this paper we will use the simple Taylor monomials $\psi_{m}(x)=x^{m}$ as basis functions. The polynomials $P_{k}^{M}(x)$ are obtained by reconstruction, i.e. by interpreting the point values $\mathbf{u}_{i,j,k}$ of the finite difference scheme as cell averages, see [72]. The associated reconstruction equations (which are not interpolation equations, see [72, 73]) for a generic polynomial of degree $M$ computed on an associated stencil $\mathcal{S}_{k}$ read
@@
    Consequently, from (50) and (51) we now compute an auxiliary polynomial $P^{N}_{0}(x)$ as
@@
    Finally, a non-linear WENO reconstruction is performed in terms of the polynomial $P^{N}_{0}(x)$ and of the lower degree polynomials built over the existing sub-stencils. In particular, we compute for all polynomials of degree $M(k)$ the nonlinear WENO oscillation indicators as usual in the following manner
@@
so that the corresponding high order accurate approximation of the first spatial derivative can be obtained as
@@ III.2 CWENO finite difference discretization of the Einstein-Euler system
The full Einstein-Euler equations (14)–(II.2) for the coupled evolution of matter and spacetime form a first–order PDE system in which the matter sector enters as a purely conservative sub–system with algebraic sources, while the Einstein sector is purely non–conservative. On the overall it can be written as
@@
where $\boldsymbol{u}$ is the state vector, composed of 63 dynamical variables [^12] More specifically: 5 for the matter part, 11 for the lapse, the shift vector, the metric components and the scalar $\phi$, 7 for $\tilde{A}_{ij}$ and the scalar $K$, 3 for $b^{i}$, 3 for $\tilde{\Gamma}^{i}$ and 1 for $K_{0}$; plus the auxiliary variables needed for the first order reduction: 3 for $A_{i}$, 9 for $B_{k}^{\,i}$, 18 for $D_{kij}$ and 3 for $P_{i}$. In comparison, the classical second order BSSNOK system [16] has only 30 evolution variables, while the well-known first order generalized harmonic (GH) formulation [13] of the full Einstein-Euler equations requires 55 variables, namely 50 for the Einstein sector and 5 variables for the matter part., while $\boldsymbol{F}=\left(\boldsymbol{f}^{x},\boldsymbol{f}^{y},\boldsymbol{f}^{z}\right)^{T}$ is the flux vector. A finite difference WENO scheme for the vacuum Einstein equations, namely with $T^{\mu\nu}=0$ and which can successfully cope with the presence of non–conservative terms, has been recently proposed by [70]. Here we follow a similar approach, with a few modifications. Still in semi-discrete form, a path-conservative finite difference scheme for the discretization of (59), reads
@@
Note that one only needs the eigenvalues, not the matrix itself. Eq. (61) and (62) should be regarded as four–state flux formulas, where each of the four entries $\boldsymbol{w}^{-}_{i+\frac{1}{2},j,k}$, $\boldsymbol{w}^{+}_{i+\frac{1}{2},j,k}$, $\boldsymbol{f}^{x,-}_{i+\frac{1}{2},j,k}$ and $\boldsymbol{f}^{x,+}_{i+\frac{1}{2},j,k}$ are reconstructed via the CWENO strategy, as described in Sect. III.1. The whole procedure must of course be repeated for the calculation of $\boldsymbol{f}^{y}_{i,j\pm\frac{1}{2},k}$ and $\boldsymbol{f}^{z}_{i,j,k\pm\frac{1}{2}}$. According to [78, 70] the three fractions involving the $\mathbf{D}^{x,y,z}$ terms in Eq. (60) are equivalent to the jump terms in path-conservative finite volume schemes [85, 86, 87, 88] and they are given as follows:
  • Rusanov–type scheme:
@@
The approximation of the terms $\tilde{\boldsymbol{B}}^{x}_{i+\frac{1}{2},j,k}$ in (67) and (III.2) is achieved through a simple midpoint rule as
@@
where $\xi_{i}^{c}$ sets the border of the inner uniform grid, while the coefficients $a$, $b$, $c$ and $d$ are selected in such a way to obtain a smooth matching of the cubic with the inner linear branch. The same mapping is applied along all coordinate directions. It is also worth stressing that, though not equivalent to AMR in terms of adaptability, the usage of a grid stretching like (72), or similar ones, reduces the need for AMR in many circumstances. An extensive usage of coordinate mapping in the context of numerical relativity has been presented in [90].
Non-reflecting boundary conditions.
Our code allows for two different strategies to obtain non-reflecting boundary conditions. Primarily, we use Sommerfeld boundary conditions for the generic quantity $Q$ as [91]
$\partial_{t}Q+\frac{r}{x_{i}}\partial_{i}Q=-\frac{1}{r}(Q-Q_{0})|_{\partial\Omega}\,,$ (73)
where $r=\sqrt{x^{2}+y^{2}+z^{2}}$, while $Q_{0}$ denotes the initial conditions. As an alternative, more pragmatic, approach, we have used a so-called sponge layer boundary condition, which performs a weighted average among the numerical solution and the initial condition at the grid border, resulting in an absorption of waves. In practice, in a thin shell at the border we modify the solution according to
$Q|_{\partial\Omega}=(1-k)Q|_{\partial\Omega}+kQ_{0}|_{\partial\Omega}\,,$ (74)
where $k$ is a small number chosen typically as $k=0.1$.
@@
For stationary problems, our numerical scheme can easily be made well–balanced, by simply subtracting a discrete version of the equilibrium solution from the discretized time-dependent PDE system. In this way, an initial equilibrium can be maintained stationary up to machine precision virtually for ever. We recall, however, that well–balancing (WB) is not a necessary requirement of our numerical scheme. It is instead an additional feature that is advantageous only for stationary solutions and it can be activated to increase the accuracy of the computation. In this paper we have used it only for the TOV star in Sect. IV.6 to highlight its potential advantages.
The interested reader is referred to [92, 93, 28], where he/she can find all the practical details for the simple and straightforward implementation of a well–balanced scheme in a PDE system of mixed conservative///non–conservative form. For more theoretical background on well-balanced numerical methods for hyperbolic PDE, see [94, 95, 96, 97, 98, 85, 86, 87] and references therein.